OpenStudy (anonymous):

how would i prove the identity (sin^2x - 6sinx+9)/sin^2x-9 = (sinx-3)/sinx+3

OpenStudy (campbell_st):

look at factorising the left hand side (sin^2(x) - 6sin(x) + 9 = (sin(x) -3)^2 sin^(x) - 9 = (sin(x)-3)(sin(x) + 3) take sin(x) -3 out as a common factor \[\frac{(\sin(x)-3)^2}{(\sin(x)-3)(\sin(x)+3)} = \frac{\sin(x) -3}{\sin(x) +3}\]

OpenStudy (anonymous):

= (sin x - 3)^2 ----------- (sinx -3 )(sinx + 3) cancelling out sinx - 3 gives the desired result

OpenStudy (anonymous):

@campbell_st @joeywhite thank you :)